### v2

# MS-TVNet:A Long-Term Time Series Prediction Method Based on Multi-Scale Dynamic Convolution

Abstract
Long-term time series prediction has predominantly relied on Transformer and MLP models, while the potential of convolutional networks in this domain remains underexplored. To address this gap, we introduce a novel multi-scale time series reshape module, which effectively captures the relationships among multi-period patches and variable dependencies. Building upon this module, we propose MS-TVNet(Multi-Scale Time-Video Network), a multi-scale 3D dynamic convolutional neural network. Through comprehensive evaluations on diverse datasets, MS-TVNet demonstrates superior performance compared to baseline models, achieving state-of-the-art (SOTA) results in long-term time series prediction. Our findings highlight the effectiveness of leveraging convolutional networks for capturing complex temporal patterns, suggesting a promising direction for future research in this field.The code is released on https://github.com/Curyyfaust/TVNet.
Index Terms:
Long-Term Time Series Forecasting, CNN, Multi-Scale, Fast Fourier Transform;Dynamic Convolution
I Introduction
Long-term time series prediction is a crucial task with extremely wide applications in professional fields such as energy management[1][2], weather forecasting[3], and health monitoring[4]. However, long-term time series forecasting tasks often involve complex multi-period scales[5][6] and cross-variable dependencies[7], making the development of effective forecasting models a significant challenge. Real-world time series exhibit diverse variations across different temporal scales. To perform accurate time series forecasting, it is essential to adopt multi-scale modeling approaches that can capture features and dependencies from various time intervals. As illustrated in Figure 1, the same time series can be divided into smaller segments (in blue) or larger segments (in yellow), resulting in fine-grained or coarse-grained temporal features. In the figure, black arrows represent relationships between adjacent time steps, capturing local details, while colored arrows indicate relationships spanning longer time steps, reflecting global dependencies[8][9][10].
Figure 1: Time series are divided into patches of varying sizes as temporal resolution[11]
Currently, convolutional models have achieved notable progress in long-term time series forecasting. For instance, SCINet[12] discards causal convolutions and employs a recursive downsampling and convolutional interaction architecture to model complex temporal dynamics. Despite these advancements in capturing long-range dependencies, significant challenges persist in long-term forecasting tasks. TimesNet[13] presents an innovative approach by transforming one-dimensional time series into two-dimensional representations and leveraging 2D convolution techniques from computer vision to enhance feature representation. ModernTCN[14] extends the application of convolution by utilizing large convolutional kernels to capture global temporal features. However, these models still fail to fully exploit the multi-scale characteristics of time series and the direct relationships across variables.
Multi-scale feature modeling has proven to be highly effective for correlation learning and feature extraction in fields such as computer vision and multimodal learning, but its potential has not been fully explored in time series forecasting. Existing methods like N-HiTS[15] model features at different resolutions through multi-resolution data sampling and hierarchical interpolation; Pyraformer[16] introduces a pyramid attention mechanism to extract features at various time resolutions; Scaleformer[17] proposes a multi-scale framework, but it increases model complexity due to the need to allocate predictive models at different time resolutions.
Through the aforementioned research, we have identified several issues in long-term time series forecasting at the current stage: (1) existing models lack adaptive multi-scale feature extraction; (2) there is a lack of a CNN-based network to extract patches and cross-variable dependencies within time series. To address these issues, this paper proposes MS-TVNet, an adaptive multi-scale 3D dynamic convolutional network for long-term time series forecasting. The innovations can be summarized as follows: (1) we proposed an adaptive Patch module based on the Fourier transform to extract multi-scale features; (2) on this basis, a multi-scale dynamic convolution is proposed to extract patches and cross-variable dependencies. We have achieved results on public datasets that surpass the baseline state-of-the-art (SOTA) models.
The organization of the article is as follows: Section II provides a detailed description of the methods; Section III presents the experiments and results; Section IV concludes the article.
II Method
II-A Problem Statement
The goal of long-term time series prediction is to leverage a given historical sequence, denoted as $\mathcal{X}_{\text{in}}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{L}]\in\mathbb{R}^{L\times C},$to accurately forecast a subsequent sequence, $\mathcal{X}_{\text{out}}=[\mathbf{x}_{L+1},\ldots,\mathbf{x}_{L+P}]\in\mathbb{R}^{P\times C}$, Where, $L$ represents the length of the historical time frame, which is the duration of the data we have observed and can use as a basis for predictions. $P$ stands for the length of the predictive time frame, indicating how far into the future aim to forecast. Meanwhile, $C$ signifies the dimensionality of the time series variables, meaning that each data point $\mathbf{x}_{i}$ in the sequence is a vector with $C$ features, capturing different aspects or features of the phenomenon being modeled.
II-B Framework
The proposed method is illustrated in Figure 2. The overall architecture of MS-TVNet is based on a Mixer-based framework and is composed of three key modules: Multi-Patch, 3D Dynamic Convolution, and Adaptive Aggregation. Each of these modules plays a distinct role in enhancing Multi-Scale feature extraction and fusion. The Multi-Patch module is designed to capture local and global temporal dependencies across multiple time scales and the 3D Dynamic Convolution module dynamically adjusts its receptive field to model complex multi-scale temporal interactions more effectively. Finally, the Adaptive Aggregation module integrates features from multiple paths to produce time series representations. In the following sections, we provide a detailed discussion of each component and its contribution to the model’s overall performance.
Figure 2: The Framework of MS-TVNet
Multi-Patch Module: This module utilizes the Fourier transform to extract multiple dominant periods embedded within the time series adaptively. By identifying these intrinsic periodic patterns, the time series can be segmented into patches that align with the discovered periods, enabling better structural representation and improved modeling efficiency. For a time series of length $L$ with $C$ features, the original format is represented as $\mathcal{X}_{\text{in}}\in\mathbb{R}^{L\times C},$ where each row corresponds to a specific timestamp and each column denotes a particular variable. To effectively capture temporal dependencies across varying scales, it is crucial first to identify the underlying periodic components. This is achieved by transforming the time series into the frequency domain using the Fast Fourier Transform (FFT), which enables the detection of frequency peaks corresponding to dominant periodicities. The resulting spectral information serves as the foundation for adaptive patch segmentation, as detailed below:
$\{f_{1},\cdots,f_{k}\}$ $=\underset{f_{*}\in\{1,\cdots,\left\lfloor\frac{T}{2}\right\rfloor\}}{\arg\text{Topk}}(\mathbf{\text{Avg}\left(\text{Amp}\left(\text{FFT}(\mathcal{X}_{\text{in}})\right)\right)})$ (1)
$p_{i}$ $=\left\lfloor\frac{T}{f_{i}}\right\rfloor,\quad i\in\{1,\cdots,k\}.$
Where $\text{FFT}(\cdot)$ and $\text{Amp}(\cdot)$ represent the Fast Fourier Transform and the calculation of amplitude values, respectively. The amplitude of each frequency is computed by averaging across $C$ dimensions using $\text{Avg}(\cdot)$. This corresponds to the patch length $p_{i}=\frac{T}{f_{i}}$, where $f_{i}$ denotes the frequency.
3D dynamic Convolution: We first embed the input time series along the feature dimension, mapping it from $C$ to a high-dimensional space of $C_{m}$, resulting in the embedded representation $\mathcal{X}_{\text{in}}\in\mathbb{R}^{L\times C_{m}}$. This step enhances the feature expressiveness. To segment the sequence into temporal patches($p_{i}$), we apply a one-dimensional convolutional layer (Conv1D) with a kernel size of $P$, effectively dividing $\mathcal{X}_{\text{in}}$ into $N=L/P$ patches. The output after convolution is denoted as $\mathcal{X}_{\text{emb}}\in\mathbb{R}^{N\times P\times C_{m}}$, where each patch preserves local temporal patterns over $P$ time steps. Next, to further capture intra-patch structure, we split each patch into two equal-length sub-patches of size $P/2$, and then stack them along a new dimension, yielding the reshaped representation $\mathcal{X}_{\text{emb}}\in\mathbb{R}^{N\times 2\times(P/2)\times C_{m}}$. This formulation enables the model to learn more nuanced dependencies within each patch. When using multiple patch lengths $p_{i}$, we can obtain a set of multi-scale embeddings $\{\mathcal{X}^{1}_{\text{emb}},\ldots,\mathcal{X}^{k}_{\text{emb}}\}$, enabling the model to capture dynamics at different temporal resolutions.
For each $\mathcal{X}^{i}_{emb}$ we propose a dynamic convolutional method to capture the temporal dynamics within each patch. To clarify, the embedded representation $\mathcal{X}^{i}_{\text{emb}}$ is envisioned as a collection of individual patches, denoted as $(x_{1},x_{2},\ldots,x_{N})$. The weight associated with the $i$-th patch, denoted as $W_{i}$, is hypothesized to be decomposable into the product of a constant time base weight $W_{b}$, which is common across all patches, and a time varying weight $\alpha_{i}$, which is unique to each patch.(see in Figure 3)
$\tilde{x}_{i}=W_{i}\cdot x_{i}=(\alpha_{i}\cdot W_{b})\cdot x_{i},$ (2)
To more accurately model the temporal dynamics of patches, the time-varying weight $\alpha_{i}$ must account for all patches. We design an adaptive time-varying weight generation that considers the interaction between inter-patch and intra-patch, articulated as $\alpha_{i}=\mathcal{G}(\mathcal{X}_{\text{emb}})$. The function $\mathcal{F}(\mathbf{v}_{\text{intra}})$ captures intra-patch features, initially applying 3D Adaptive Average Pooling to obtain $\mathbf{v}_{\text{intra}}$. For inter-patch channel modeling, a single-layer Conv1D is applied. The equations are as follows:
$\alpha_{i}=\mathcal{G}(\mathcal{X}_{\text{emb}})=1+\mathcal{F}(\mathbf{v}_{\text{intra}})+\mathcal{F}(\mathbf{v}_{\text{inter}})$ (3a)
$\mathbf{v}_{\text{intra}}=\text{AdaptiveAvgPool3d}(\mathcal{X}_{\text{emb}})$ (3b)
$\mathcal{F}_{\text{intra}}(\mathbf{v}_{\text{intra}})=\delta(\text{BN}(\text{Conv1D}^{C\rightarrow C}(\mathbf{v}_{\text{intra}})))$ (3c)
$\mathbf{v}_{\text{inter}}=\text{AdaptiveAvgPool1d}(\mathbf{v}_{\text{intra}})$ (3d)
$\mathcal{F}_{\text{inter}}(\mathbf{v}_{\text{inter}})=\delta(\text{Conv1D}^{C\rightarrow C}(\mathbf{v}_{\text{inter}}))$ (3e)
Here, $\delta$ denotes the ReLU activation function.
Figure 3: The Time varying weight generation flow chart
After extracting features from each $\mathcal{X}^{i}_{\text{emb}}$ using dynamic convolution, we obtain the 3D feature vector set $\{\mathcal{\hat{X}}^{1}_{\text{emb}},\ldots,\mathcal{\hat{X}}^{k}_{\text{emb}}\}$.
Adaptive aggregation:involves fusing $k$ distinct 3D-feature vectors $\{\mathcal{\hat{X}}^{1}_{\text{emb}},\ldots,\mathcal{\hat{X}}^{k}_{\text{emb}}\}$ for the subsequent layer. The amplitudes $\mathcal{A}$ reflect the relative importance of selected frequencies and periods, corresponding to the importance of each transformed 3D tensor. We aggregate the 1D-representations based on these amplitudes:
$\mathcal{A}_{f_{1}},\cdots,\mathcal{A}_{f_{k}}$ $=\arg\text{Topk}\left(\text{Avg}\left(\text{Amp}\left(\text{FFT}(\mathcal{X}_{\text{in}})\right)\right)\right)$ (4a)
$\hat{\mathcal{A}}_{f_{1}},\cdots,\hat{\mathcal{A}}_{f_{k}}$ $=\text{Softmax}\left(\mathcal{A}_{f_{1}},\cdots,\mathcal{A}_{f_{k}}\right)$ (4b)
$\mathcal{X}_{\text{emb}}$ $=\sum_{i=1}^{k}\hat{\mathcal{A}}_{f_{i}}\times\mathcal{\hat{X}}^{i}_{\text{emb}}$ (4c)
The final prediction results are obtained using residual connections and an MLP, as shown in the following formulas:
$Y=MLP(\mathcal{X}_{emb}+\mathcal{X}_{in})$ (5)
III Experiment and Result
In this section, we evaluate the performance of MS-TVNet on public datasets using Mean Squared Error (MSE) and Mean Absolute Error (MAE) as our metrics. It is worth noting that, for these metrics, a lower value indicates a superior performance. To fully validate the effectiveness of the proposed method in this paper, we conducted comparisons on six public datasets (Exchange, ETTm2, ETTh1, Electricity, Weather, ILI) with nine baselines( TVNet[7], PatchTST[18], iTransformer[19], Crossformer[20], RLinear[21], MTS-Mixer[22], DLinear, TimesNet[13], and MICN[23]).
Our method and baselines are optimized using the Adam optimizer[24]. The learning rate starts at $1e^{-4}$ and is halved when loss stagnates. The loss function used is MSE. We evaluate accuracy using $MSE=\frac{1}{n}\sum_{i=1}^{n}(y-\hat{y})^{2}$ and $MAE=\frac{1}{n}\sum_{i=1}^{n}|y-\hat{y}|$. All models are run on a single NVIDIA 3090 GPU with 24 GB of memory.
TABLE I: Forecasting result comparison with different horizon lengths. The lookback length is set to 24 for ILI and 96 for the others. Bold indicates the best result, while underlining indicates the second-best result.
Models MS-TVNet TVNet PatchTST iTransformer Crossformer RLinear MTS-Mixer DLinear TimesNet MICN
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Exchange 96 0.087 0.217 0.080 0.195 0.093 0.214 0.086 0.206 0.186 0.346 0.083 0.301 0.083 0.200 0.081 0.203 0.107 0.234 0.102 0.235
192 0.159 0.291 0.163 0.285 0.192 0.312 0.177 0.299 0.467 0.522 0.170 0.293 0.174 0.296 0.157 0.293 0.226 0.344 0.172 0.316
336 0.284 0.392 0.291 0.394 0.350 0.432 0.331 0.417 0.783 0.721 0.309 0.401 0.336 0.417 0.305 0.414 0.367 0.448 0.272 0.407
720 0.642 0.601 0.658 0.594 0.911 0.716 0.847 0.691 1.367 0.943 0.817 0.680 0.900 0.715 0.643 0.601 0.964 0.746 0.714 0.658
ETTm2 96 0.137 0.250 0.161 0.254 0.165 0.255 0.180 0.264 0.421 0.461 0.164 0.253 0.177 0.259 0.167 0.260 0.187 0.267 0.178 0.273
192 0.204 0.281 0.220 0.293 0.220 0.292 0.250 0.309 0.503 0.519 0.219 0.290 0.241 0.303 0.224 0.303 0.249 0.309 0.245 0.316
336 0.264 0.318 0.272 0.316 0.274 0.329 0.311 0.348 0.611 0.580 0.273 0.326 0.297 0.338 0.281 0.342 0.312 0.351 0.295 0.350
720 0.347 0.365 0.349 0.379 0.362 0.385 0.412 0.407 0.996 0.750 0.366 0.385 0.396 0.398 0.397 0.421 0.497 0.403 0.389 0.406
ETTh1 96 0.354 0.392 0.371 0.408 0.370 0.399 0.386 0.405 0.386 0.429 0.366 0.391 0.372 0.395 0.375 0.399 0.384 0.402 0.396 0.427
192 0.390 0.404 0.398 0.409 0.413 0.421 0.441 0.436 0.419 0.444 0.404 0.412 0.416 0.426 0.405 0.416 0.557 0.436 0.430 0.453
336 0.402 0.408 0.401 0.409 0.422 0.436 0.487 0.458 0.440 0.461 0.420 0.423 0.455 0.449 0.439 0.443 0.491 0.469 0.433 0.458
720 0.437 0.446 0.458 0.459 0.447 0.466 0.503 0.491 0.519 0.524 0.442 0.456 0.475 0.472 0.472 0.490 0.521 0.500 0.474 0.508
Electricity 96 0.139 0.220 0.142 0.223 0.129 0.222 0.148 0.240 0.187 0.283 0.140 0.235 0.141 0.243 0.153 0.237 0.168 0.272 0.159 0.267
192 0.160 0.238 0.165 0.241 0.147 0.240 0.162 0.253 0.258 0.330 0.154 0.248 0.163 0.261 0.152 0.249 0.184 0.289 0.168 0.279
336 0.159 0.260 0.164 0.269 0.163 0.259 0.178 0.269 0.323 0.369 0.171 0.264 0.176 0.277 0.169 0.267 0.198 0.308 0.161 0.259
720 0.176 0.273 0.190 0.284 0.197 0.290 0.225 0.317 0.404 0.423 0.209 0.297 0.212 0.308 0.233 0.344 0.220 0.320 0.203 0.312
Weather 96 0.140 0.198 0.147 0.198 0.149 0.198 0.174 0.214 0.153 0.217 0.175 0.225 0.156 0.206 0.152 0.237 0.172 0.220 0.161 0.226
192 0.186 0.235 0.194 0.238 0.194 0.241 0.221 0.254 0.197 0.269 0.218 0.260 0.199 0.248 0.220 0.282 0.219 0.261 0.220 0.283
336 0.230 0.269 0.235 0.277 0.245 0.282 0.278 0.296 0.252 0.311 0.265 0.294 0.249 0.291 0.265 0.319 0.280 0.306 0.275 0.328
720 0.304 0.324 0.308 0.331 0.314 0.334 0.358 0.347 0.318 0.363 0.329 0.339 0.336 0.343 0.323 0.362 0.365 0.359 0.311 0.356
ILI 24 1.319 0.705 1.324 0.712 1.319 0.754 2.207 1.032 3.040 1.186 4.337 1.507 1.472 0.798 2.215 1.081 2.317 0.934 2.684 1.112
36 1.182 0.765 1.190 0.772 1.430 0.834 1.934 0.951 3.356 1.230 4.205 1.481 1.435 0.745 1.963 0.963 1.972 0.920 2.507 1.013
48 1.396 0.695 1.456 0.782 1.553 0.815 2.127 1.004 3.441 1.223 4.257 1.484 1.474 0.822 2.130 1.024 2.238 0.940 2.423 1.012
60 1.624 0.782 1.652 0.796 1.470 0.788 2.298 0.998 3.608 1.302 4.278 1.487 1.839 0.912 2.368 1.096 2.027 0.928 2.653 1.085
1st Count 37 5 4 0 0 0 0 1 0 1
The forecasting results are comprehensively presented in Table I. A clear trend is observed where the prediction accuracy tends to decline progressively as the forecast horizon length increases, reflecting the inherent challenges in long-term forecasting. In the comparative analysis of various forecasting models, MS-TVNet consistently outperforms its counterparts by achieving the MSE and MAE across multiple datasets and forecast horizons. This consistent performance underscores its robustness and ability to maintain high accuracy even in more complex forecasting scenarios. The visualization of these prediction results is further illustrated in Figure 4, which offers a clear graphical representation of the model’s predictive capacity.
Figure 4: The Forecasting result for ETTh1($L$=96))
Computational Efficiency:Figure 5 illustrates the efficiency and memory usage of different models. It can be seen from the figure that Transformer-based models have higher runtime (s/epoch) and memory usage compared to CNN-based models. Our proposed method does not employ the stacking approach used in TVNet, resulting in lower memory usage and runtime compared to TVNet.This indicates that the method proposed in this paper can achieve a balance between efficiency and utility.
Figure 5: Model efficiency comparison under the setting of L(prediction length) =192 of ETTm2
Input Length: We evaluated the performance of our proposed method across a range of input lengths. As illustrated in Figure 6, the prediction performance of the Transformer-based model deteriorates as the input length increases. This may be attributed to the fact that the attention mechanism in Transformer-based models tends to capture periodic information incorrectly as the look-back window lengthens. In contrast, the method proposed in this paper fully accounts for periodicity. Therefore, as the look-back window expands, more periodic information can be captured, leading to a reduction in prediction error.
Figure 6: The MSE results with different input lengths and same prediction lengths
IV Conclusion
In conclusion, MS-TVNet, our proposed multi-scale 3D dynamic convolutional neural network, has shown superior performance in long-term time series forecasting. With its innovative modules—such as Multi-Patch, 3D Dynamic Convolution, and Adaptive Aggregation—MS-TVNet effectively captures intricate temporal patterns and dependencies, offering high accuracy in predictions. Comprehensive evaluations across various datasets confirm its state-of-the-art performance, establishing its potential as a powerful tool for time series forecasting. These findings highlight the promising capabilities of convolutional networks in this domain, while suggesting potential avenues for future research, including architecture enhancements and domain-specific adaptations.